### v1

# ERA-IT: Aligning Semantic Models with Revealed Economic Preference for Real-Time and Explainable Patent Valuation

Abstract
Valuing intangible assets under uncertainty remains a critical challenge in the strategic management of technological innovation due to the information asymmetry inherent in high-dimensional technical specifications. Traditional bibliometric indicators, such as citation counts, fail to address this friction in a timely manner due to the systemic latency inherent in data accumulation. To bridge this gap, this study proposes the Economic Reasoning Alignment via Instruction Tuning (ERA-IT) framework. We theoretically conceptualize patent renewal history as a revealed economic preference and leverage it as an objective supervisory signal to align the generative reasoning of Large Language Models (LLMs) with market realities, a process we term Eco-Semantic Alignment. Using a randomly sampled dataset of 10,000 European Patent Office patents across diverse technological domains, we trained the model not only to predict value tiers but also to reverse-engineer the Economic Chain-of-Thought from unstructured text. Empirical results demonstrate that ERA-IT significantly outperforms both conventional econometric models and zero-shot LLMs in predictive accuracy. More importantly, by generating explicit, logically grounded rationales for valuation, the framework serves as a transparent cognitive scaffold for decision-makers, reducing the opacity of black-box AI in high-stakes intellectual property management.
1 Introduction
In the contemporary knowledge-driven economy, patents, as core intangible assets, are increasingly pivotal to firms’ strategic decisions and value creation ernst2003 . They serve as the fundamental currency of innovation, delineating a firm’s technological trajectory and securing market dominance hall2005 ; reitzig2004 . Accurate and timely patent valuation forms the bedrock of firms’ strategy for managing technological innovation  fischer2014 ; kalip2022 . However, achieving objective and accurate ex-ante valuation of patents remains a classic and persistent grand challenge in information science, information economics, and technological innovation management reitzig2004 . Figure 1 outlines the evolutionary path of patent valuation methodologies.
Figure 1: Evolutionary trajectory of patent valuation methodologies. The figure illustrates the paradigm shift from subjective manual assessment and latency-prone bibliometric indicators to the proposed ERA-IT framework, which enables real-time, explainable valuation via generative reasoning.
The crux of this challenge lies in two facts. First, patents themselves are complex and information-dense hybrid "technical-legal" artifacts. Their true value is deeply encoded in unstructured technical texts han2015 . This creates profound information asymmetry between informed patent holders and external evaluators—such as investors and managers—who struggle to penetrate this “technical black box” and to transact efficiently in markets for technology akerlof1970 ; gambardella2007 ; gans2008 . This opacity creates severe market friction, where stakeholders struggle to distinguish high-value inventions from the vast noise of global filings, due to the high-dimensional complexity of mapping technical specifications to market outcomes pakes1986 ; trajtenberg1990 . Second, information asymmetry is inherent in the global patent ecosystem. Unlike domain-specific studies that benefit from homogeneous terminologies, a generalized valuation framework must contend with heterogeneous linguistic patterns across disparate technological sectors pakes1984 . In this context, traditional bibliometric indicators, such as citation counts, fail to serve as effective signals for timely decision-making due to systemic temporal latency; a patent typically accumulates meaningful citation data only years after its publication harhoff2003 . This "backward-looking" nature of bibliometrics prevents managers from assessing patent value in a way that is both real-time and closely reflects genuine market preferences. Consequently, there exists an urgent imperative to bridge the gap between technical disclosure and economic realization. This imperative drives the quest for advanced architectures capable of leveraging "forward-looking" signals, constituting a critical research frontier.
In this context, historical data on patent renewal has gained theoretical prominence as a direct signal of "revealed economic preferences", since the act of renewal represents a "vote with the wallet" that is more closely aligned with market reality than any external proxy  pakes1986 ; pakes1989 ; schankerman1986 . From the perspective of economic signaling theory, the decision to pay monotonically increasing renewal fees reveals the patent holder’s preference, indicating that the asset’s private value exceeds its maintenance expenses kabore2012 ; kumar2024 . Unlike stochastic citations, renewal records provide a definitive, discrete signal of value maintenance available throughout the patent’s lifecycle. However, a fundamental methodological gap persists: how to effectively align this structured, low-dimensional economic-behavior signal with the high-dimensional, unstructured semantic information in patent texts to build an accurate and explainable predictive model. Existing frameworks have struggled to integrate this objective economic signal with the rich, unstructured semantic information embedded within patent documents arts2018 . Traditional Natural Language Processing (NLP) techniques often fail to capture the nuanced causal relationships between the specific exclusionary scope defined in the claims and the patent’s potential for economic longevity lee2020 .
The recent advent of Large Language Models (LLMs) offers a transformative opportunity to overcome these linguistic barriers. LLMs possess the unique capacity for complex reasoning and for extracting latent semantic features. Nevertheless, applying general-purpose LLMs to patent valuation is non-trivial; these models often lack the domain-specific intuition required to align technical novelty with economic viability. To overcome this critical barrier in high-stakes Intellectual Property (IP) management in such contexts, there is an urgent need for adaptive technological and methodological innovations. The development of a predictive architecture that effectively fuses the objective ground truth of renewal data with the semantic reasoning capabilities of LLMs, a process that we term "Eco-Semantic Alignment”, represents a largely unexplored frontier in intelligent information systems zhang2022 .
Figure 2: Conceptual overview of the ERA-IT framework. The architecture integrates high-dimensional semantic context processed by LLMs with objective economic signals derived from renewal data to generate explainable valuation outcomes.
To bridge this methodological gap, this study proposes the Economic-Reasoning Alignment via Instruction Tuning (ERA-IT) framework, as illustrated in Figure 2. Our approach uses European Patent Office (EPO) renewal data as a supervisory signal to guide LLMs’ generative reasoning. Guided by a rigorous data fusion strategy, we operationalize the longitudinal payment of renewal fees as a value proxy. This allows us to translate economic behavior into a labeled dataset with discrete value tiers. This transforms the ambiguous task of valuation into a robust supervised learning problem. Subsequently, we employ a strategy that synergizes parameter-efficient fine-tuning with chain-of-thought prompting. This alignment enables the model to learn latent semantic patterns correlated with economic survival, effectively converting unstructured text into quantifiable predictive indicators.
Specifically, we explore three critical research questions underexplored in prior studies:
  • RQ1 (Efficacy): Does ERA-IT outperform state-of-the-art discriminative baselines and zero-shot LLMs in predicting patent value?
  • RQ2 (Mechanism): How does the explicit generation of economic reasoning contribute to the model’s predictive performance?
  • RQ3 (Generalizability): Is the framework robust across heterogeneous technological domains (International Patent Classification (IPC) sections) and low-resource settings?
To rigorously validate the efficacy and generalizability of this framework, we conducted a comprehensive empirical analysis using a dataset comprising over 10,000 EPO patents, randomly sampled across various IPC sections. This randomized sampling strategy ensures that our findings are not biased toward specific high-velocity domains and instead represent a robust solution applicable to the broader technological landscape. Our results demonstrate that this hybrid approach yields significantly superior predictive accuracy compared to conventional rule-based heuristics and statistical baselines arts2018 . By minimizing predictive latency in citation metrics, our system provides a scalable solution for automated high-precision IP valuation.
Our study pioneers a theoretical framework by aligning LLMs’ semantic reasoning with revealed economic preferences to establish a new paradigm for real-time and interpretable patent valuation. It has at least threefold research contributions, summarized as follows.
  • Theoretical Contribution via Eco-Semantic Alignment: We introduce a novel theoretical framework that aligns high-dimensional semantic representations with objective economic renewal signals. By integrating economic signaling theory with natural language processing, this approach reveals how discrete market signals can effectively guide the decoding of complex, unstructured technical specifications, resolving the label ambiguity inherent in unsupervised text analysis. We demonstrate how, by designing specific alignment mechanisms (i.e., Eco-Semantic Alignment), general-purpose AI technology (LLMs) can be successfully adapted to the complex context of IP valuation. This task is highly dependent on professional knowledge, and this adaptation provides a new pathway for realizing intelligent, precise, and reliable IP public services and the strategic management of technological innovation.
  • Methodological Contribution in Causal Reasoning: We establish the algorithmic efficacy of the ERA-IT architecture, which synergizes parameter-efficient fine-tuning with domain-specific instruction tuning strategies. We provide empirical evidence that standard pre-trained models are insufficient for the specialized task of IP valuation, demonstrating that our domain-adaptation process captures the "Economic Chain-of-Thought" within obfuscated legal-technical texts.
  • Generalizability and Managerial Implications: We validate the model’s capacity to extract universal value signals through extensive empirical analysis of a randomly sampled dataset. Unlike prior studies that were restricted to specific sectors, this work demonstrates the model’s applicability across heterogeneous patent texts. By mitigating the systemic latency of bibliometric metrics, the proposed framework empowers stakeholders to make real-time data-driven decisions on the deployment of R&D resources and the commercialization of technology.
Overall, our study provides a novel interdisciplinary lens that bridges information science, information economics, technological innovation management, and explainable AI, thereby enriching the theoretical foundation of context-adaptive innovation. Coherently, it makes contributions through the theoretical conception of Eco-Semantic Alignment and its methodological instantiation, the ERA-IT framework. This synergy offers a new paradigm for the deep adaptation of AI in professional domains, one that is anchored in objective economic reality and realized through an implementable, verifiable algorithmic framework with strong generalization capabilities and rigorous evidentiary standards. As a pioneering study, this work achieves AI innovation in a specific professional context by advancing patent valuation from a field reliant on lagging proxy indicators to a new stage driven by real-time market data and equipped with an interpretable reasoning process.
2 Related Works
2.1 Patent Valuation
Research on patent valuation has shifted from simple patent counts toward richer indicators that proxy the economic and innovative impact of inventions. Using the American Inventors Protection Act, which mandated earlier publication of most US patent applications, Hegde et al. (2023) show that earlier disclosure is associated with more and faster follow-on citations, fewer abandonments, and higher R&D investment, highlighting how publication rules shape both the private and social value of patents. Svensson (2022) analyzes four widely used patent value indicators, renewal length, forward citations, family size, and documents that each capture distinct facets of technological and economic value rather than a single underlying measure. Building on this indicator-based tradition, Ponta et al. (2021) propose the Innovation Patent Index (IPI), a composite innovation-performance measure that aggregates multiple patent characteristics related to technological quality, market potential, and scope into a firm-level index. Complementing these firm- and patent-level perspectives, Park et al. (2023) apply the previously introduced CD index of disruption to large-scale datasets of papers and patents, extend it with normalized variants and alternative measures, and show that science and technology have become less disruptive over time, illustrating the potential of advanced citation-based metrics for characterizing the qualitative role of inventions in the knowledge network.
However, from the perspective of ex-ante patent valuation, these approaches remain largely ex-post and rely on outcome variables that materialize only after a substantial time lag. Renewal behavior, composite indices such as IPI, and disruption-based citation metrics are typically calibrated on long-term renewal patterns, realized commercialization, or accumulated citation structures, and they make limited use of the full semantic content of patent texts at the level of individual inventions Svensson (2022); Ponta et al. (2021); Park et al. (2023). Moreover, these indicators generally summarize value at the firm or field level and do not provide explicit, stepwise economic rationales for why a particular patent should be assessed as more or less valuable at the time of filing. Because of these limitations, existing patent-valuation approaches remain constrained in their ability to provide economically grounded and explainable ex-ante value assessments for newly filed patents. These limitations underscore the need for ex-ante patent valuation frameworks that integrate semantic text analysis with interpretable reasoning, such as ERA-IT.
2.2 Semantic Analysis for IP Management
Recent work in IP management increasingly relies on semantic analysis of patent and related technical text. Ali et al. (2024) review the evolution of patent retrieval and group methods into families that range from keyword and vector space models to semantic and deep learning-based techniques for prior art search. Xu and Ono (2025) extend this line of work to industry–academia collaboration by proposing a semantic retrieval approach that uses text embeddings to analyze and rank collaborative research outputs. At the model level, Chikkamath et al. (2022) construct large-scale USPTO subclass datasets and investigate how a BERT for Patents encoder and strong non-neural baselines use claims and abstracts for multi-label patent classification. Complementary research on structured representations decomposes Chinese patent claims into function, structure, and purpose segments and applies pre-trained transformer-based models to classify them into technical categories, thereby improving classification accuracy and supporting retrieval of patents in similar technical domains Li et al. (2025). In parallel, Zini and Awad (2022) survey explainability techniques for deep NLP models, cataloguing model-agnostic and model-specific methods that interpret embeddings, internal representations, and output decisions for text-based tasks.
Despite these advances, existing semantic approaches are optimised for retrieval, classification, or document alignment rather than for transparent ex-ante economic valuation of patents. The patent retrieval survey focuses on relevance and recall, leaving the relationship between semantic similarity and economic value largely implicit Ali et al. (2024). Semantic matching systems for industry–academia collaboration operate in high-dimensional embedding spaces but do not attempt to infer monetary or strategic value or to incorporate ground-truth financial signals Xu and Ono (2025). Models such as the BERT for Patents and the structured claim classifiers in the Chinese patent study are trained to predict technology codes, not renewal-based value tiers, and typically output numerical labels without explicit economic rationales Chikkamath et al. (2022); Li et al. (2025). Even where explainability is the primary topic, current surveys emphasise post-hoc interpretation tools for generic NLP tasks and note that rigorous evaluation in high-stakes legal and economic contexts remains limited Zini and Awad (2022). In contrast, ERA-IT explicitly fuses semantic reasoning with revealed preference renewal data and uses instruction-tuned chain-of-thought generation to map claim language to value tiers, positioning ERA-IT as a bridge between semantic IP analytics and transparent economic valuation while addressing both the opacity of black-box language models and the latency of traditional citation-based indicators.
2.3 Generative AI for IP Management
Recent work on instruction-tuned generative models provides the technical basis for structured reasoning in ERA-IT. Instruction tuning surveys indicate that LLMs can be further trained on collections of instruction-response pairs to improve controllability and domain adaptation while systematising common pitfalls of instruction-driven alignment zhang2023 . Chain-of-Thought prompting studies demonstrate that demonstrations with explicit intermediate steps can markedly improve multi-step reasoning, and that relevance and step ordering matter more than strictly correct logical derivations wang2023cot . Research on self-explanations finds that model-generated rationales may appear plausible yet diverge from traditional attribution methods, motivating explicit checks of explanation faithfulness huang2023selfexplain . In high-stakes legal settings, participatory design with practising lawyers argues that LLMs should primarily function as decision-support tools within explicit policy frameworks rather than as autonomous legal advisers cheong2024law . Within the IP domain, domain-specific frameworks guide models along patent-relevant dimensions. For example, PatentMind uses a multi-aspect reasoning graph to decompose and score patent similarity across technical, legal, and contextual dimensions yoo2025patentmind ; PatentScore defines a multi-dimensional rubric for evaluating LLM-generated claims across structural, semantic, and legal aspects yoo2025patentscore ; and Self Filtered Distillation uses LLM-generated labels and rationales to compute trust scores, thereby filtering or reweighting training examples for patent classification yoo2025sfd .
However, from the perspective of ex-ante patent valuation, these techniques remain only loosely connected to economic ground truth. Instruction tuning and Chain-of-Thought prompting are typically optimised on general reasoning benchmarks and do not directly constrain rationales or predictions with observable economic outcomes like renewal behaviour, so their explanations are at best indirectly related to value signals zhang2023 ; wang2023cot . Studies of self-explanations further show that natural language rationales can appear faithful under coarse metrics while assigning very different importance patterns, which complicates their use as evidence for economic reasoning huang2023selfexplain . In the patent domain, PatentMind and PatentScore align model behaviour with expert assessments of similarity and claim quality, and Self Filtered Distillation improves the reliability of patent classifiers using trust indicators derived from model outputs, yet none of these frameworks explicitly tie reasoning to renewal trajectories or other revealed preference signals yoo2025patentmind ; yoo2025patentscore ; yoo2025sfd . Therefore, ERA-IT couples patent renewal history, treated as revealed economic preference, with a four-step Economic Chain-of-Thought schema covering technical strength, market context, legal robustness, and value synthesis, and uses teacher-guided data synthesis together with parameter-efficient adaptation to align LLM reasoning with these economic signals and to turn generated explanations into actionable features for ex-ante patent valuation.
3 Dataset Construction
3.1 Sampling Strategy and Longitudinal Alignment
To establish a valuation framework with high external validity, it is critical to mitigate the selection bias often present in domain-specific studies. Unlike prior works restricted to high-velocity sectors, such as semiconductors or biotech, we aimed to capture the universal linguistic patterns of economic value applicable across the entire patent ecosystem. Accordingly, our data collection followed a rigorous, randomized, stratified sampling strategy. We targeted historical patent filings within the EPO database, specifically censoring the observation window to the 2010–2015 cohort. This timeframe ensures an ex post observation horizon of at least eight years, allowing for a complete retrospective analysis of economic survival trajectories. The final analytical corpus comprises 10,000 patent documents uniformly sampled across all primary IPC sections (A–H). By enforcing this heterogeneity, we ensure that the model learns to extract generalized economic signals rather than overfitting to the terminologies of specific industries. Table 1 presents representative examples derived from the constructed ground-truth dataset.
Table 1: Representative examples of the constructed ground-truth dataset. The corpus consists of 10,000 patents filed between 2010 and 2015, randomly sampled across diverse IPC sections (A-H). This randomized stratification ensures that the proposed framework captures universal value signals across heterogeneous technological landscapes, minimizing domain-specific overfitting.
ID IPC Section Filing Year Subject Matter (Summary) Renewals Assigned Label
Patent 1 A (Human Necessities) 2011 Manual coffee brewing mechanism 2 yrs Class 1 (Low)
Patent 2 B (Transporting) 2013 Vehicle side mirror housing assembly 5 yrs Class 2 (Mid)
Patent 3 G (Physics/Computing) 2014 Neural network quantization method 9 yrs Class 3 (High)
Patent 4 C (Chemistry) 2012 CRISPR-Cas9 gene editing vector 11 yrs Class 3 (High)
Patent 5 H (Electricity) 2010 Legacy 3G communication protocol 3 yrs Class 1 (Low)
3.2 Operationalizing Value via Revealed Preferences
To construct objective ground-truth labels for supervised learning, we operationalized patent renewal records as a proxy for private economic value, grounded in the theory of revealed preferences kabore2012 ; harhoff2003 . Theoretically, the payment of annual renewal fees constitutes a "costly signal". It reflects the patent holder’s internal assessment that the expected future returns (or option value) of the asset exceed the cumulative and monotonically increasing maintenance expenses pakes1986 . In the context of information systems, this financial decision serves as a definitive binary signal at each time step, encoding the latent utility of the invention free from the noise of "cheap talk". Unlike citation metrics, which suffer from significant lag in accumulation and strategic manipulation, renewal behavior provides a real-time, forward-looking indicator of value retention. This allows us to convert discrete financial transactions into a robust target variable for algorithmic prediction kumar2024 .
3.3 Label Discretization and Value Stratification
To convert the continuous-time series of renewal events into a format suitable for classification, the corpus was stratified into three discrete value tiers. This discretization aligns with the typical hazard rate distribution of patent longevity observed in the econometric literature kumar2024 ; kabore2012 .
  • Class 1 (Low Value - Early Screening): Patents renewed one to three times. This tier represents the lower quartile of the value distribution, characterizing inventions where rights were relinquished during the initial screening phase. In economic terms, this rapid abandonment signals that the patent holder realized the technology’s lack of commercial viability or technical obsolescence shortly after the grant.
  • Class 2 (Medium Value - Standard Cycle): Patents renewed four to six times. These patents exhibit a standard technological lifespan, reflecting moderate economic relevance. They represent the median behavior of the dataset, where the asset generates sufficient utility to justify maintenance costs during the initial commercialization phase but lacks the strategic importance to warrant long-term protection against escalating fees.
  • Class 3 (High Value - Strategic Assets): Patents renewed seven or more times. This category captures the "long tail" of high-value IP. These patents were maintained for an extended period despite the EPO’s typically high fee structure. Continued renewal at this stage serves as a strong signal of sustained competitive advantage, strategic blocking value, or significant revenue generation capability.
4 Methodology: ERA-IT Framework
In this section, we propose the ERA-IT framework. Unlike traditional discriminative approaches that treat patent valuation as a black-box classification task, our framework adopts a generative paradigm to model the causal reasoning process underlying economic valuation explicitly. By aligning LLMs’ semantic processing capabilities with objective market signals, we aim to bridge the gap between high-dimensional technical specifications and their realized private value. Figure 3 illustrates the overall pipeline of the proposed architecture.
Figure 3: The schematic overview of the ERA-IT framework. The pipeline integrates teacher-guided data synthesis with parameter-efficient instruction tuning to align the model with economic reasoning.
4.1 Problem Formulation: From Classification to Causal Reasoning
Traditional patent valuation is typically formulated as maximizing the conditional probability $P(y|X)$, where $y\in\mathcal{Y}$ represents the discrete value tier given the patent document $X$. However, this formulation suffers from interpretative opacity; it fails to explicate "why" a patent holds value, often relying on spurious correlations within the text.
We redefine this task as a Conditional Reasoning Generation problem, aiming to model the Economic Chain-of-Thought. Our objective is to generate a sequence $S=(r,y)$ that includes $r$, the logical rationale justifying the valuation (e.g., interpretation of exclusionary scope, analysis of barriers to entry), followed by the prediction $y$. Consequently, the model parameters $\theta$ are optimized to maximize the following likelihood:
$\theta^{*}=\operatorname*{argmax}_{\theta}\sum_{(X,r,y)\in\mathcal{D}}\log P_{\theta}(r,y|X)$ (1)
where $\mathcal{D}$ represents the instruction tuning dataset constructed to capture the causal link between technical disclosure and economic survival.
4.2 Reverse-Engineering the Economic Chain-of-Thought
A critical bottleneck in supervised learning for valuation is the scarcity of high-quality reasoning data ($r$). To address this, we propose a semi-automated pipeline that uses patent renewal data as objective ground truth to reverse-engineer the latent economic logic.
4.2.1 Structural Schema for Cognitive Alignment
To ensure the generated reasoning aligns with domain experts’ cognitive processes, we define a four-step reasoning schema as follows. This schema forces the model to sequentially evaluate the structural components of patent value before reaching a conclusion.
  • Technical Identification ($\phi_{tech}$): The model distills the technical essence of the invention, filtering out obfuscated legalese to identify the core novelty.
  • Market Application Analysis ($\phi_{market}$): The model infers the potential industrial application and market size, estimating the demand-side dynamics.
  • Exclusionary Scope Interpretation ($\phi_{legal}$): The model evaluates the Appropriability of the invention by analyzing the breadth of independent claims and the difficulty of design-around by competitors.
  • Valuation Synthesis ($\phi_{value}$): Synthesizing the technical, market, and legal factors, the model establishes a causal link to the realized renewal duration.
4.2.2 Teacher-Guided Data Synthesis
For the training corpus $X_{train}$ with actual renewal history, we leverage a high-performance Teacher Model (specifically, GPT-4) to synthesize the reasoning data $\hat{r}$. Figure 4 illustrates the process of using the teacher model to conduct data synthesis. We employ a Label-Conditioned Generation strategy by providing the ground-truth label $y_{gt}$ as an anchor in the prompt. This induces the model to construct an ex post rationalization based on the actual economic outcome. The generated $\hat{r}$ serves as the pseudo-cognitive path that logically bridges the document $X$ and the value $y$. The final dataset is curated as $\mathcal{D}_{CoT}=\{(X_{i},r_{i},y_{i})\}_{i=1}^{N}$.
Figure 4: The data synthesis process utilizing a Teacher Model to reverse-engineer economic reasoning based on ground-truth renewal data.
4.3 Domain Adaptation via Instruction Tuning
We fine-tune a Student LLM (Llama-3-8B) using the constructed $\mathcal{D}_{CoT}$. To effectively inject domain-specific economic logic while preserving the model’s general linguistic capabilities, we adopt a parameter-efficient training strategy using LoRA (Low-Rank Adaptation).
4.3.1 Instruction-Formatted Input Representation
The raw patent document $X$ is transformed into a structured prompt in hierarchical tag format. The input sequence $I$ is defined as follows:
$I=\text{[SYS]}\oplus\mathcal{I}_{task}\oplus\text{[META]}\oplus M\oplus\text{[DOC]}\oplus X\oplus\text{[ANS]}$ (2)
where $\mathcal{I}_{task}$ denotes the role-playing instruction (e.g., "You are an expert in IP analyst"), forcing the model to adopt the persona of a valuation expert.
4.3.2 LoRA-based Adaptation
We freeze the pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$ and update the learning parameters by introducing low-rank approximation matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$, where the rank is $r\ll\min(d,k)$. The weight update is formulated as:
$W=W_{0}+\Delta W=W_{0}+\frac{\alpha}{r}BA$ (3)
This approach allows for securing domain alignment by training less than 0.1% of the parameters, preventing catastrophic forgetting of general reasoning capabilities.
4.4 Optimization Objective with Instruction Masking
Model training follows the Causal Language Modeling loss function. Crucially, we apply Instruction Masking to prevent the model from merely memorizing the input patent text $X$.
Let $T=[t_{1},\dots,t_{L}]$ be the tokenized sequence of the formatted input $I$, and $L_{input}$ be the length of the prompt segment (instruction and document). The optimization loss function $\mathcal{L}$ is defined as:
$\mathcal{L}(\theta)=-\sum_{i=L_{input}+1}^{L}\log P_{\theta}(t_{i}|t_{1},\dots,t_{i-1};X)$ (4)
By calculating the loss exclusively on the economic rationale and value tier segments ($r$ and $y$), we force the model to focus on learning the conditional probability distribution that yields a logical valuation consistent with the specification. This distinguishes our framework from traditional approaches that predict outcomes without an explicit reasoning process.
4.5 Inference and Value Prediction
During the inference phase, the model is provided with the input prompt $I_{test}$ containing only the instruction and the patent document. The model generates the token sequence autoregressively:
$S_{gen}=\operatorname*{argmax}_{S}P_{\theta}(S|I_{test})$ (5)
Since the model is trained to generate the output in a structured format, we employ a deterministic parsing function $\psi(\cdot)$ to extract the predicted value tier $\hat{y}$. This end-to-end generation process allows stakeholders to obtain not only the valuation score but also the interpretative transparency required for cognitive scaffolding in decision-making.
5 Experiments
To rigorously validate the proposed ERA-IT framework, we designed comprehensive experiments guided by the three research questions proposed in Section 1.
5.1 Experimental Setup
5.1.1 Datasets and Evaluation Metrics
We utilized the dataset of 10,000 EPO patents constructed in Section 4.2. The dataset was stratified and randomly split into training (8,000), validation (1,000), and testing (1,000) sets. Given the potential class imbalance in real-world scenarios, we report not only Accuracy but also Macro-F1 score and the Matthews Correlation Coefficient (MCC), which is considered a more robust metric for imbalanced classification tasks.
5.1.2 Baselines
We compared ERA-IT against three categories of strong baselines:
  • Traditional machine learning: TF-IDF+SVM and TF-IDF+RF serve as fundamental benchmarks for text classification.
  • Discriminative pre-trained language models: We included BERT-Large, SciBERT (trained on scientific text), and Longformer (optimized for long documents) to evaluate the limitations of encoder-only models.
  • Generative LLMs: We evaluated Deepseek-r1-8B, Llama-3-8B, Qwen-3-8B (base model) in a 5-shot setting and GPT-5-mini (zero-shot) to measure the gap between general reasoning and domain-specific instruction tuning.
5.1.3 Implementation Details
The Student AI (Llama-3-8B) was fine-tuned using LoRA with rank $r=16$, $\alpha=32$, and dropout $0.05$. We used the AdamW optimizer with a cosine learning rate schedule (initial lr=2e4lr=2e-4) and a warm-up ratio of 0.03. All experiments were conducted on a single NVIDIA H100 GPU, ensuring reproducibility.
5.2 Performance Analysis (RQ1)
The ERA-IT framework achieves superior performance across all metrics, as shown in Table 2. Notably, it outperforms Longformer, suggesting that generative reasoning is more effective for valuation than merely processing long contexts. Furthermore, ERA-IT surpasses GPT-5-mini, demonstrating that parameter-efficient fine-tuning with economic signals yields better domain alignment than the general reasoning capability of closed-source giant models.
Table 2: Comparative performance on the test set.† indicates statistical significance (p < 0.01) compared to the best baseline.
Model Type Model Accuracy Macro-F1 MCC
Traditional TF-IDF + SVM 58.4 56.2 0.38
TF-IDF + Random Forest 61.2 59.8 0.42
Discriminative BERT-Large 69.5 68.1 0.54
SciBERT 73.2 71.8 0.59
Longformer 74.8 73.2 0.62
Generative Deepseek-r1-8b (5-shot) 64.2 61.8 0.51
Llama-3-8B (5-shot) 68.9 67.4 0.52
Qwen-3-8B (5-shot) 69.4 69.1 0.49
GPT-5-mini (Zero-shot) 76.1 73.5 0.65
Ours ERA-IT (Full) 83.4 79.6 0.79
5.3 Ablation and Mechanism Analysis (RQ2)
To isolate the contribution of the Economic Chain-of-Thought reasoning, we conducted an ablation study, as shown in Table 3.
Table 3: Ablation study on model components. The removal of each component leads to a performance drop, confirming their contributions to the framework.
Configuration Macro-F1 ($\Delta$)
ERA-IT (Full Framework) 79.6
   w/o Rationale Generation ($r$) 74.5 (-5.1)
   w/o Instruction Masking 77.2 (-2.4)
   w/o Hierarchical Tags (Flat Text) 78.4 (-1.2)
The most significant performance drop (-5.5%) occurs when the rationale generation ($r$) is removed. This confirms that the generated reasoning acts as a "cognitive scaffold", forcing the model to attend to relevant technical and legal details before making a prediction. Removing instruction masking also degrades performance, indicating that focusing the loss solely on the output helps the model learn the generation task more efficiently.
5.4 Robustness and Generalizability (RQ3)
5.4.1 Performance Across Technology Domains
To verify the generalizability claimed in the Introduction, we analyzed the model’s accuracy across different IPC sections. As showm in Table 5, the model maintains high accuracy ($80\%$) in Physics (G) and Electricity (H), and shows robust performance even in Chemistry (C) and Mechanics (B), confirming its applicability across heterogeneous patent landscapes.
Figure 5: Prediction accuracy across major IPC sections. The model demonstrates consistent performance regardless of the technological domain.
5.4.2 Sample Efficiency
We further evaluated the model’s performance under low-resource settings by varying the training data size from 10% to 100%. ERA-IT achieves 90% of its peak performance with only 40% of the data, suggesting that the Economic Chain-of-Thought data provides dense supervisory signals, enabling efficient learning even with limited labeled data.
5.5 Qualitative Case Study
[…]
Table 4: A generated inference example for a Class 3 patent (Solid-state Electrolyte).
Input Claim Fragment:
[…]
Model Generated Rationale ($r$):
1. Technical Identification: The invention specifies a solid-state electrolyte composition, focusing on particle distribution for safety.
[…]
3. Market Viability: This feature is critical for EV battery safety, implying high commercial demand.
Predicted Value Tier ($y$): Class 3 (High Value)
6 Conclusion
6.1 Summary of Research Findings
In this study, we addressed the persistent challenges of information asymmetry and interpretative opacity in the patent market by proposing the ERA-IT framework. Moving beyond traditional bibliometric approaches that suffer from significant temporal latency, we theoretically redefined patent renewal history as a revealed economic preference. This enabled us to establish a novel supervised learning paradigm that leverages objective market signals to guide the semantic decoding of LLMs. Our empirical analysis, conducted on a dataset of 10,000 EPO patents randomly sampled across diverse technological domains, substantiates the efficacy of this approach. The ERA-IT framework not only achieved a significant performance gain over state-of-the-art discriminative baselines (Macro-F1 +7.2%) and zero-shot LLMs but also demonstrated robust generalizability across heterogeneous IPC sections ranging from Chemistry to Physics. Crucially, the model successfully reverse-engineers the Economic Chain-of-Thought, providing logical rationales that bridge the epistemic gap between obfuscated technical specifications and their realized market value.
6.2 Theoretical Implications
This research offers substantial theoretical contributions to the literature on information systems, AI innovation, and IP Management by establishing Eco-Semantic Alignment as a new paradigm connecting economic reality with AI semantic reasoning.
First, we introduce revealed economic preference as a novel supervision benchmark for the AI domain adaptation. This study points out that traditional citation-based bibliometric indicators suffer from systemic latency and cannot provide effective signals for real-time valuation. In contrast, the patent renewal decision is a market action taken by the right holder based on private information, directly reflecting the patent’s expected economic value. The core theoretical contribution of this research lies in systematically demonstrating that patent renewal history—a discrete, observable market behavior—serves as an objective and timely ground truth for large language models in understanding the semantics of patent texts. This finding challenges the prevailing reliance on citation-based indicators, proving that financial maintenance decisions offer a more immediate and less noisy proxy for technological value. By directly mapping linguistic patterns to economic survival, our work transcends the label ambiguity inherent in traditional unsupervised text analysis. It provides a clear, stable anchor in economic logic for decoding complex technical documents, thereby expanding the theoretical understanding of how value is encoded within technical documentation. By algorithmically merging high-dimensional, unstructured technical text with low-dimensional yet information-rich economic behavior data, the ERA-IT framework improves the efficiency and precision with which firms convert technical information into strategic intelligence. This theoretically empowers firms to conduct context-adaptive innovation in an increasingly digital, intelligent, and dynamic environment.
Second, we define the Economic Chain-of-Thought as the mechanism to bridge the cognitive gap between semantic representation and market value, extending the application of Generative AI from simple text summarization to Causal Reasoning in Economic Contexts. Our framework does not stop at the black-box prediction of "input text, output value." Its key theoretical innovation is proposing the concept of the Economic Chain-of-Thought and making it an explainable output that the model must generate. This concept theoretically aligns the semantic reasoning process of AI with the economic causal logic of human decision-makers. It requires the model not only to predict an outcome but also to reverse-engineer, from obscure legal-technical texts, the chain of reasoning that supports a specific economic decision (to renew or not). This essentially builds a transparent cognitive scaffold for the AI’s reasoning, making its outputs no longer opaque but rational arguments that decision-makers can review, question, and adopt. By showing that instruction tuning with domain-specific economic logic significantly outperforms standard pre-training, this study establishes a new benchmark for "Vertical AI" applications where domain expertise is scarce, high-dimensional, and requires logical transparency.
Third, we establish a computable theoretical pathway for firms to achieve AI’s context-adaptive innovation. This study provides a concrete instantiation of this concept: by using a domain-specific, objective economic behavior signal (renewal) as the supervision signal for instruction tuning, it compels a general-purpose LLM to adapt its internal semantic representations to the logic of the specific context of IP valuation. This alignment is not superficial keyword matching but drives the model’s generative reasoning to conform to the behavioral logic of micro-economic agents, thereby achieving a cognitive-level adaptation of the AI system from "general-purpose" to "domain expert".
Overall, our study builds a new paradigm to resolve a core contradiction in the field: the chasm between technical complexity and economic measurability.
6.3 Methodological Implications
This research designs and validates the ERA-IT executable architecture for achieving Eco-Semantic Alignment. First, we develop a hybrid training strategy that co-optimizes parameter-efficient fine-tuning and domain-specific instruction. Addressing the dual challenges of patent text specialization and the scarcity of high-quality labeled data, the ERA-IT method does not rely on computationally intensive full-parameter fine-tuning. Its methodological innovation lies in the synergistic integration of parameter-efficient fine-tuning techniques with a carefully constructed set of domain instructions that guide economic reasoning. This approach explicitly shows that for tasks requiring deep domain reasoning, such as patent valuation, standard pre-trained models are insufficient. Instead, a strategy combining lightweight parameter adaptation with explicit task instructions can more effectively capture domain-specific logic while avoiding overfitting.
Second, we provide causal empirical evidence for the establishment of the Economic Chain-of-Thought to model performance. The empirical component of the research goes beyond simply demonstrating improved predictive accuracy. Through structured experimental designs (e.g., ablation studies), it provides causally inferable evidence that the model’s significant performance gain is attributable to its ability to generate the Economic Chain-of-Thought. This methodologically confirms that the performance gain does not stem from memorizing superficial text features but from the model successfully learning and simulating the internal logical chain connecting technical semantics to economic decisions. This sets a new standard for evaluating the effectiveness of AI models in professional domains: superior performance must be accompanied by an auditable reasoning process that aligns with domain logic.
Third, we validate the method’s robust cross-technological domain generalization using Economically Universal Signals. Another key methodological contribution is demonstrating the framework’s robust generalizability through testing across different IPC sections. This benefits from the nature of its supervision signal (renewal behavior)—it is a universal economic logic that transcends specific technical details. This indicates that anchoring on such high-level, cross-domain universal behavioral signals is an effective methodological strategy for building specialized AI systems capable of handling technological diversity and overcoming data sparsity challenges in niche domains.
6.4 Managerial and Practical Implications
This study holds significant managerial and practical value. The proposed framework serves as a transformative decision-support system for R&D strategists, patent attorneys, and venture investors, specifically addressing the cost of due diligence.
  • Mitigating Information Asymmetry via Real-time Valuation: Innovation markets suffer from severe information asymmetry. Our proposed ERA-IT framework addresses this fundamental challenge by introducing patent renewal history as a revealed preference signal to align LLMs’ generative reasoning with market realities, thereby bridging the gap between informed patent holders and external evaluators, such as investors and managers. Unlike citation metrics, which require years to accumulate, our model enables stakeholders to assess the potential value of patent applications immediately upon publication. This enables agile portfolio management and timely technology-transfer decisions in high-velocity markets. This framework enables firms to conduct real-time, automated economic health scans of their entire patent portfolio. Managers can quickly identify "dormant value" patents with a strong economic rationale and relatively low maintenance costs, as well as "non-performing assets" with a weak economic rationale but that incur high renewal fees.
  • Cognitive Scaffolding and Human-AI Complementarity: The ability of ERA-IT to generate explicit rationales acts as a cognitive scaffold for human experts. Instead of replacing human judgment, we propose a "Propose-and-Verify" workflow to ensure high-trust collaboration. First, the AI functions as a high-throughput screening agent that generates value predictions alongside a transparent Economic Chain-of-Thought. Second, the domain expert engages in "Rapid Verification" by reviewing the generated rationale to confirm its legal and technical validity, rather than analyzing the raw specification from scratch. This process leverages the complementarity between the computational scalability of AI and the strategic intuition of human experts, thereby minimizing the black-box risk associated with deep learning deployment. Third, the Economic Chain-of-Thought is not intended to replace IP managers, technical experts, or lawyers but to serve as their "cognitive augmentation partner". It provides a structured basis for discussion and a common language upon which human experts can conduct reviews, questionings, refinements, and final decision-making. This helps to externalize and institutionalize the tacit knowledge of top experts, while improving the overall rationality and consistency of team decisions.
In summary, by grounding LLMs’ semantic reasoning in the objective economic reality of the market, this study offers a powerful new approach to valuing and managing intangible assets. It provides both a telescope for strategic foresight and a microscope for real-time, transparent analysis. In doing so, our work charts a course toward a future in which IP management evolves into a precise, data-driven decision science, built on human-machine collaboration and deeply woven into the fabric of corporate strategy.
6.5 Limitations and Future Directions
Despite its significant contributions, this study is subject to limitations that delineate avenues for future research. First, while patent renewal data serves as a robust proxy for private value, it is not free from noise. Firms may maintain low-value patents for strategic blocking purposes or abandon high-value patents due to financial constraints. Future research could refine the ground-truth labels by incorporating auxiliary signals, such as litigation records or standard-essential patent declarations, to distinguish between intrinsic technological value and strategic value. Second, the "rationales" generated by our model, while logically consistent, rely on the patterns learned from the teacher model. To ensure these rationales fully align with human expert intuition, future studies should conduct rigorous human evaluations involving patent attorneys to quantify the validity and legal accuracy of the generated reasoning. Finally, our empirical validation was restricted to European Patent Office data and textual modalities. Extending the framework to other jurisdictions (e.g., the USPTO and the CNIPA) and integrating visual encoders for technical drawings would represent a significant step toward a truly global, multimodal valuation system.